# Interpretable factorization of clinical questionnaires to identify latent factors of psychopathology

### 12 Dec 2023 v1

@@ Abstract
Psychiatry research seeks to understand the manifestations of psychopathology in behavior, as measured in questionnaire data, by identifying a small number of latent factors that explain them. While factor analysis is the traditional tool for this purpose, the resulting factors may not be interpretable, and may also be subject to confounding variables. Moreover, missing data are common, and explicit imputation is often required. To overcome these limitations, we introduce interpretability constrained questionnaire factorization (ICQF), a non-negative matrix factorization method with regularization tailored for questionnaire data. Our method aims to promote factor interpretability and solution stability. We provide an optimization procedure with theoretical convergence guarantees, and an automated procedure to detect latent dimensionality accurately. We validate these procedures using realistic synthetic data. We demonstrate the effectiveness of our method in a widely used general-purpose questionnaire, in two independent datasets (the Healthy Brain Network and Adolescent Brain Cognitive Development studies). Specifically, we show that ICQF improves interpretability, as defined by domain experts, while preserving diagnostic information across a range of disorders, and outperforms competing methods for smaller dataset sizes. This suggests that the regularization in our method matches domain characteristics. The python implementation for ICQF is available at https://github.com/jefferykclam/ICQF.
Keywords:
non-negative matrix factorization questionnaire interpretability psychopathology.
@@ 1 Introduction
Standardized questionnaires are a common tool in psychiatric practice and research, for purposes ranging from screening to diagnosis or quantification of severity. A typical questionnaire comprises questions – usually referred to as items – reflecting the degree to which particular symptoms or behavioural issues are present in study participants. Items are chosen as evidence for the presence of latent constructs giving rise to the psychiatric problems observed. For many common disorders, there is a practical consensus on constructs. If so, a questionnaire may be organized so that subsets of the items can be added up to yield a subscale score quantifying the presence of their respective construct. Otherwise, the goal may be to discover constructs through factor analysis.
The factor analysis (FA) of a questionnaire matrix (#$\mathrm{\#participants\times\#items}$) expresses it as the product of a factor matrix (#$\mathrm{\#participants\times\#factors}$) and a loading matrix (#$\mathrm{\#factors\times\#items}$). The method assumes that answers to items should be correlated, and can therefore be explained in terms of a smaller number of factors. The method yields two real-valued matrices, with uncorrelated columns in the factor matrix. The number of factors is specified a priori, or estimated from data. The values of the factors for each participant can then be viewed as a succinct representation of them.
Interpreting what construct a factor may represent is done by considering its loadings across items. Ideally, if very few items have a non-zero loading, or each item only has a high loading on a single factor, it will be easy to associate the factor with them. The FA solution is often subjected to rotation to try to accomplish this. In practice, the loadings could be an arbitrary linear combination of items, with positive and negative weights. Factors are real-valued, and neither their magnitude nor their sign are intrinsically meaningful. Beyond this, any missing data will have to be imputed, or the respective items omitted, before FA can be used. Finally, patterns in answers that are driven by other characteristics of participants (e.g. age or sex) are absorbed into factors themselves, acting as confounders, instead of being represented separately or controlled for.
@@
  • Factor loadings are bounded within the same range as the original questionnaire responses, facilitating interpretation as answer patterns associated with the factor, rather than arbitrary values.
  • The reconstructed matrix adheres to the range or observed maximum of the original questionnaire, preventing any entry from exceeding these limits.
@@
2. Theoretical foundations of ICQF  Introducing constraints on both the factors and the reconstructed matrix poses algorithmic challenges. We introduce an optimization procedure for ICQF, using alternating minimization with ADMM, and we demonstrate that it converges to a local minimum of the optimization problem. We implement blockwise-cross-validation (BCV) to determine the number of factors. We show that, if this number of factors is close to that underlying the data, the solution will be close to a global minimum. We also empirically demonstrate that BCV detects the number of factors more precisely than competing methods through synthetic questionnaire examples.
@@
4. Light-weighted implementation  We provide a Python implementation of ICQF that can efficiently handle typical questionnaire datasets in psychology or psychiatry research contexts.
@@ 2 Related Work and Technical Motivation for our Method
The extraction of latent variables (a.k.a. factors) from matrix data is often done through low rank matrix factorizations, such as singular value decomposition (SVD), principal component analysis (PCA) and exploratory Factor Analysis (hereafter, just FA) [26, 8]. While SVD and PCA aim at reconstructing the data, FA aims at explaining correlations between (questions) items through latent factors [3]. Factor rotation [12, 51, 52] is then performed to obtain a sparser solution which is easier to interpret and analyze. For a review of FA, see [55, 24, 28, 27]. Non-negative matrix factorization (NMF) was proposed as a way of identifying sparser, more interpretable latent variables, which can be added to reconstruct the data matrix. It was introduced in [46] and developed in [39]. Different varieties of NMF-based models have been proposed for various applications, such as the sparsity-controlled [19, 49], manifold-regularized [41], orthogonal [16, 14], convex/semi-convex [17], or archetypal regularized NMF [35]. More recently, Deep-NMF [56, 63] and Deep-MF [62, 21, 2] can model non-linearities on top of (non-negative) factors, when the sample is large [20]. These methods do not directly model either the interpretability characteristics or the constraints that we view as desirable. If the goal is to identify latent variables relevant for multiple matrices, the standard approach is multi-view learning [54], or variants that can handle only partial overlap in participants across matrices [18, 29, 25]. Finally, non-negative matrix tri-factorization [40, 47], supports an additional matrix mapping between latent representations for different matrices.
@@
Obtaining a factorization with these methods requires both specifying the number of latent variables, and solving an optimization problem. In SVD/PCA, the number of variables is often selected based on the percentage of variance explained, or determined via techniques such as spectral analysis, the Laplace-PCA method, or Velicer’s MAP test [57, 58, 44]. For FA, several methods have been proposed: Bartlett’s test [4], parallel analysis [33, 30], MAP test and comparison data [50]. For NMF, iterative detection algorithms are recommended, e.g. the Bayesian information criterion (BIC) [53], cophenetic correlation coefficient (CCC) [22] and the dispersion [13]. More recent proposals for NMF are Bi-cross-validation (BiCV) [45] and its generalization, the blockwise-cross-validation (BCV) [36], which we use in this paper. The optimization problem for NMF is non-convex, and different algorithms for solving it have been proposed. Multiplicative update (MU) [39] is the simplest and mostly used. Projected gradient algorithms such as the block coordinate descent [15, 60, 38] and the alternating optimization [37, 42] aim at scalability and efficiency in larger matrices. Given that our optimization problem has various constraints, we use a combination of alternative optimization and Alternating Direction Method of Multipliers (ADMM) [11, 34].
@@ 3.1 Interpretable Constrained Questionnaire Factorization (ICQF)
Inputs  Our method operates on a questionnaire data matrix $M\in\mathbb{R}_{\geq 0}^{n\times m}$ with $n$ participants and $m$ questions, where entry $(i,j)$ is the answer given by participant $i$ to question $j$. As questionnaires often have missing data, we also have a mask matrix $\mathcal{M}\in\{0,1\}^{n\times m}$ of the same dimensionality as $M$, indicating whether each entry is available $(=1)$ or not $(=0)$. Optionally, we may have a confounder matrix $C\in\mathbb{R}_{\geq 0}^{n\times c}$, encoding $c$ known variables for each participant that could account for correlations across questions (e.g. age or sex). If the $j^{th}$ confound $C_{[:,j]}$ is categorical, we convert it to indicator columns for each value. If it is continuous, we first rescale it into $[0,1]$ (where 0 and 1 are the minimum and maximum in the dataset), and replace it with two new columns, $C_{[:,j]}$ and $1-C_{[:,j]}$. This mirroring procedure ensures that both directions of the confounding variables are considered (e.g. answer patterns more common the younger or the older the participants are). Lastly, we incorporate a vector of ones into $C$ to facilitate intercept modeling of dataset wide answer patterns.
@@
subject to entries of $Q$ being in the same value range as question answers, so loadings are interpretable, and bounding the reconstruction by the range of values in the questionnaire matrix $M$. We further regularize $W$ and $Q$ through $R(W,Q):=\|W\|_{p,q}+\gamma\|Q\|_{p,q}$, $\gamma=\frac{n}{m}\max(M)$, where $\|A\|_{p,q}:=(\sum^{m}_{i=1}(\sum^{n}_{j=1}|A_{ij}|^{p})^{q/p})^{1/q}$. Here, we use $p=q=1$ for sparsity control. The heuristic $\gamma$ balances the sparsity control between $W$ and $Q$; $\gamma$ is absorbed into $\beta$ of $Q$ if no ambiguity results.
@@ 3.2 Solving the optimization problem
We use the ADMM framework for fitting the ICQF model, due to its parallelizability, flexibility in incorporating various types of constraints, and its compatibility with different optimization schemes. Specifically, we utilize the Fast Iterative Shrinkage Thresholding Algorithm (FISTA) to accommodate our sparsity constraints, leveraging its numerical advantages, such as quadratic convergence and low memory cost, as discussed in [23]. Unlike stochastic optimization approaches, which require addressing the missing entries and uneven distribution of responses in questionnaires when generating training batches, ADMM allows us to tackle the optimization problem holistically. Additionally, it can find a solution for large clinical questionnaire datasets (thousands of participants, tens to hundreds of questions) in about a minute with a laptop CPU, so the performance is appropriate. Alternatively, we also provide an optimization procedure based on coordinate descent algorithm for small data sizes in the Appendix 6.2.
@@
$\begin{split}&\mathcal{L}_{\rho}(W,Q,Z,\alpha_{Z})\\ =&1/2\|\mathcal{M}\odot(M-Z)\|_{F}^{2}+\mathcal{I}_{\mathcal{W}}(W)+\beta\|W\|_{1,1}+\mathcal{I}_{\mathcal{Q}}(Q)+\beta\|Q\|_{1,1}\\ &+\left\langle\alpha_{Z},Z-[W,C]Q^{T}\right\rangle+\rho/2\left\|Z-[W,C]Q^{T}\right\|^{2}_{F}+\mathcal{I}_{\mathcal{Z}}(Z)\end{split}$
@@
Equations 1 and 2 can be further split into row-wise constrained Lasso problems and there is a closed form solution for equation 3. The optimization details are further discussed in Appendix 6.1. Given the flexibility of ADMM, a similar procedure can also be used with other regularizations.
Convergence of the optimization procedure  The convergence hinges on the careful selection of the penalty parameter $\rho$. Informally, imposing the constraint $\rho\geq\sqrt{2}$ on the penalty parameter $\rho$ guarantees monotonicity of the optimization procedure, and that it will converge to a local minimum. Integrating this constraint with the adaptive selection of $\rho$ [61], we obtain an efficient optimization procedure for ICQF. Formally, this can be stated as the following proposition.
Proposition 1 (Non-increasing property)
Assume $\rho\geq\sqrt{2}$, we have
@@
The main idea of the proof of 1 is to estimate the difference between the two consecutive Lagrangians in Equation 5 by expanding it into
$\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1)},\alpha_{Z}^{(i+1)})-\mathcal{L}_{\rho}(\mathbbm{V}^{(i)},\alpha_{Z}^{(i)})=$ $\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1)},\alpha_{Z}^{(i+1)})-\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1)},\alpha_{Z}^{(i)})$
$+\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1)},\alpha_{Z}^{(i)})-\mathcal{L}_{\rho}(\mathbbm{V}^{(i)},\alpha_{Z}^{(i)})$ (6)
where $\mathbbm{V}^{(i)}:=\left\{W^{(i)},Q^{(i)},Z^{(i)}\right\}$. Given that the subproblems 13 are minimized during each iteration, we can estimate upper bounds of these terms and obtain
@@
$\leq$ $\left(\frac{1}{\rho}-\frac{\rho}{2}\right)\cdot\left(\|[W^{(i+1)},C](Q^{(i+1),T}-Q^{(i),T})\|_{F}^{2}\right.$
$+\left.\|[(W^{(i+1)}-W^{(i)}),C]Q^{(i),T}\|_{F}^{2}+\|Z^{(i+1)}-Z^{(i)}\|_{F}^{2}\right).$ (7)
If we set $\rho\geq\sqrt{2}$, the right hand side becomes negative and the Lagrangian decreases across iterations and converges to a critical point. The full proof of Proposition 1 is given in Appendix 6.3.
@@
Furthermore, [9] showed that, for non-negative matrix factorizations, if the dimensionality $k$ is the same as that $k^{*}$ of a ground truth solution $(W^{*},Q^{*})$, the error $\|M-WQ^{T}\|_{F}^{2}$ is star-convex towards $(W^{*},Q^{*})$, and the solution is close to a global minimum. However, if $k\neq k^{*}$, the relative error between $W^{*}$ and $W$ increases with $|\sqrt{k/k^{*}}-1|$. Inaccurate estimation of $k^{*}$ thus affects both the interpretability of ($W$, $Q$) and the convergence to global minima. With the bounded constraints imposed on $W$ and $Q$ in ICQF, Popoviciu’s inequality establishes an upper bound for the variances $\sigma_{W}^{2}$ and $\sigma_{Q}^{2}$ of each column in $W$ and $Q$ respectively. To simplify the analysis, we assume equal variances among the columns (generally true). Then we have the following proposition:
Proposition 2
@@
Let $(W^{*},Q^{*})$ be a ground-truth factorization of the given $\mathbf{M}=\mathbf{W}^{*}(\mathbf{Q}^{*})^{T}$, with latent dimension $k^{*}$, where $\mathbf{W}^{*}$ and $\mathbf{Q}^{*}$ are matrix-valued random variables with entries sampled from bounded distributions. Suppose $(\mathbf{W},\mathbf{Q})$ is another factorization with dimension $k\neq k^{*}$, then

### 17 Mar 2026 v2

@@ Abstract
Psychiatry research seeks to understand the manifestations of psychopathology in behavior, as measured in questionnaire data, by identifying a small number of latent factors that explain them. While factor analysis is the canonical tool for this purpose, the resulting factors may not be interpretable, and may also be subject to confounding variables. Moreover, missing data are common, and explicit imputation is often required. To overcome these limitations, we introduce Interpretability Constrained Questionnaire Factorization (ICQF), a non-negative matrix factorization method with regularization tailored for questionnaire data. Our method aims to promote factor interpretability and solution stability. We provide an optimization procedure with theoretical convergence guarantees, and an automated procedure to determine latent dimensionality accurately. We validate these procedures using realistic synthetic data. We demonstrate the effectiveness of our method in a widely used general-purpose questionnaire, in two independent datasets (the Healthy Brain Network and Adolescent Brain Cognitive Development studies). Specifically, we show that ICQF improves interpretability, as defined by domain experts, while preserving diagnostic information across a range of disorders, and outperforms competing methods for smaller dataset sizes. This suggests that the regularization in our method matches domain characteristics.
@@ 1 Introduction
Standardized questionnaires are a common tool in psychiatric practice and research, for purposes ranging from screening to diagnosis or quantification of severity. A typical questionnaire comprises questions – usually referred to as items – reflecting the degree to which particular symptoms or behavioural issues are present in study participants. The items provide evidence for the presence of latent constructs giving rise to the psychiatric problems observed. For many common disorders, there is a practical consensus on constructs. If so, a questionnaire may be organized so that subsets of the items can be added up to yield a subscale score quantifying the presence of their respective construct. Otherwise, the goal may be to discover new constructs, by factorizing a questionnaire containing items of interest in a particular domain.
The factor analysis (FA) of a questionnaire matrix ($\mathrm{\#participants\times\#items}$) expresses it as the product of a factor matrix ($\mathrm{\#participants\times\#factors}$) and a loading matrix ($\mathrm{\#factors\times\#items}$). The method assumes that answers to items should be correlated, and can therefore be explained in terms of a smaller number of factors. The method yields two real-valued matrices, with uncorrelated columns in the factor matrix. The number of factors is specified a priori, or estimated from data. The values of the factors for each participant can then be viewed as a succinct representation of them.
Interpreting what construct a factor may represent is done by considering its loadings across items. Ideally, if very few items have a non-zero loading, or each item only has a high loading on a single factor, it will be easy to associate the factor with them. The FA solution is often subjected to rotation to accomplish this. In practice, the loadings could be an arbitrary linear combination of items, with positive and negative weights. Factors are real-valued, and neither their magnitude nor their sign are intrinsically meaningful. Beyond this, any missing data will have to be imputed, or the respective items omitted, before FA can be used. Finally, patterns in answers that are driven by other characteristics of participants (e.g. age or sex) are absorbed into factors themselves, acting as confounders, instead of being represented separately or controlled for.
@@
  • Factor loadings are bounded within the same value range as the original questionnaire responses, facilitating interpretation as answer patterns associated with the factor, rather than arbitrary values.
  • The reconstructed matrix adheres to the known value range or observed maximum of the original questionnaire, preventing any entry from exceeding these limits.
@@
2. Theoretical foundations of ICQF  Introducing constraints on both the factors and the reconstructed matrix poses algorithmic challenges. We introduce an optimization procedure for ICQF, using alternating minimization with ADMM, and we demonstrate that it converges to a local minimum of the optimization problem. We implement blockwise-cross-validation (BCV) to determine the number of factors. We show that, if this number of factors is close to that underlying the data, the solution will be close to a global minimum. We also empirically demonstrate that BCV detects the number of factors more precisely than competing methods, through synthetic questionnaire examples.
@@
4. Light-weighted implementation  We provide a Python implementation of ICQF that can efficiently handle questionnaire datasets in psychology or psychiatry research contexts, with large-scale participant cohorts.
@@ 2 Related Work and Technical Motivation for our Method
The extraction of latent variables (a.k.a. factors) from matrix data is often done through low rank matrix factorizations, such as singular value decomposition (SVD), principal component analysis (PCA) and exploratory Factor Analysis (FA) Golub & Van Loan (2013); Bishop & Nasrabadi (2006). While SVD and PCA aim at reconstructing the data, FA aims at explaining correlations between (questions) items through latent factors Bandalos & Boehm-Kaufman (2010). Factor rotation Browne (2001); Sass & Schmitt (2010); Schmitt & Sass (2011) is then performed to obtain a sparser solution which is easier to interpret and analyze. For a review of FA, see Thompson (2004); Gaskin & Happell (2014); Gorsuch (2014); Goretzko et al. (2021). Non-negative matrix factorization (NMF) was proposed as a way of identifying sparser, more interpretable latent variables, which can be added to reconstruct the data matrix. It was introduced in Paatero & Tapper (1994) and developed in Lee & Seung (2000). Different varieties of NMF-based models have been proposed for various applications, such as the sparsity-controlled Eggert & Korner (2004); Qian et al. (2011), manifold-regularized Lu et al. (2012), orthogonal Ding et al. (2006); Choi (2008), convex/semi-convex Ding et al. (2008), or archetypal regularized NMF Javadi & Montanari (2020). More recently, Deep-NMF Trigeorgis et al. (2016); Zhao et al. (2017) and Deep-MF Xue et al. (2017); Fan & Cheng (2018); Arora et al. (2019) can model non-linearities on top of (non-negative) factors, when the sample is large Fan (2021). These methods do not directly model either the interpretability characteristics or the constraints that we view as desirable.
@@
Obtaining a factorization with these methods requires first specifying the number of latent variables, and then solving an optimization problem. In SVD/PCA, the number of variables is often selected based on the percentage of variance explained, or determined via techniques such as spectral analysis, the Laplace-PCA method, or Velicer’s MAP test Velicer (1976); Velicer et al. (2000); Minka (2000). For FA, several methods have been proposed: Bartlett’s test Bartlett (1950), parallel analysis Horn (1965); Hayton et al. (2004), MAP test and comparison data Ruscio & Roche (2012). For NMF, iterative detection algorithms are recommended, e.g. the Bayesian information criterion (BIC) Stoica & Selen (2004), cophenetic correlation coefficient (CCC) Fogel et al. (2007) and the dispersion Brunet et al. (2004). More recent proposals for NMF are Bi-cross-validation (BiCV) Owen & Perry (2009) and its generalization, the blockwise-cross-validation (BCV) Kanagal & Sindhwani (2010), which we use in this paper. The optimization problem for NMF is non-convex, and different algorithms for solving it have been proposed. Multiplicative update (MU) Lee & Seung (2000) is the simplest and mostly used. Projected gradient algorithms such as the block coordinate descent Cichocki & Phan (2009); Xu & Yin (2013); Kim et al. (2014) and the alternating optimization Kim & Park (2008); Mairal et al. (2010) aim at scalability and efficiency in larger matrices. Given that our optimization problem has various constraints, we use a combination of alternative optimization and Alternating Direction Method of Multipliers (ADMM) Boyd et al. (2011); Huang et al. (2016).
@@ 3.1 Interpretable Constrained Questionnaire Factorization (ICQF)
Inputs  Our method operates on a questionnaire data matrix $M\in\mathbb{R}_{\geq 0}^{n\times m}$ with $n$ participants and $m$ questions, where entry $(i,j)$ is the answer given by participant $i$ to question $j$. As questionnaires often have missing data, we also have a mask matrix $\mathcal{M}\in\{0,1\}^{n\times m}$ of the same dimensionality as $M$, indicating whether each entry is available $(=1)$ or not $(=0)$. Optionally, we may have a confounder matrix $C\in\mathbb{R}_{\geq 0}^{n\times c}$, encoding $c$ known variables for each participant that could account for correlations across questions (e.g. age or sex). If the $j^{th}$ confound $C_{[:,j]}$ is categorical, we convert it to indicator columns for each value. If it is continuous, we first rescale it into $[0,1]$ (where 0 and 1 are the minimum and maximum in the dataset, or known range), and replace it with two new columns, $C_{[:,j]}$ and $1-C_{[:,j]}$. This mirroring procedure ensures that monotonic effects in either direction of the confounding variables are representable (e.g. answer patterns more common the younger or the older the participants are), while preserving non-negativity. Importantly $C$ is not used to regress out confounds in the classical linear regression sense. Instead, it acts as a non-negative baseline (offset) components that captures systematic response patterns attributable to known variables. Under the non-negativity constraints, this confound contribution models additive baseline structure, allowing the latent factors to explain residual variation rather than compensate for it. Lastly, we incorporate a vector of ones into $C$ to model dataset-wide intercept effects, to capture capture global answer tendencies across the population.
@@
subject to entries of $Q$ being in the same value range as question answers, so loadings are interpretable, and bounding the reconstruction by the range of values in the questionnaire matrix $M$. We regularize $W$ and $Q$ through $R(W,Q):=\|W\|_{p,q}+\gamma\|Q\|_{p,q}$ with $\gamma=\frac{n}{m}\max(M)$, where $\|A\|_{p,q}:=(\sum^{m}_{i=1}(\sum^{n}_{j=1}|A_{ij}|^{p})^{q/p})^{1/q}$. Here, we use $p=q=1$ to promote sparsity. The scaling factor $\gamma$ normalizes the regularization across factors with different dimensions and value ranges ($W\in[0,1]$ versus $Q\in[\min(M),\max(M)]$), preventing disproportionate penalization of either factor and mitigating the intrinsic scaling ambiguity of matrix factorization, thereby improving numerical stability. Accordingly, we use a single regularization parameter $\beta$ to control the overall sparsity level, treating $\gamma$ as a fixed normalization constant absorbed into the penalty of $Q$.
@@ 3.2 Solving the optimization problem
We use the ADMM framework for fitting the ICQF model, due to its parallelizability, flexibility in incorporating various types of constraints, and its compatibility with different optimization schemes. Specifically, we utilize the Fast Iterative Shrinkage Thresholding Algorithm (FISTA) to accommodate our sparsity constraints, leveraging its numerical advantages, such as quadratic convergence and low memory cost, as discussed in Gaines et al. (2018). Unlike stochastic optimization approaches, which require addressing the missing entries and uneven distribution of responses in questionnaires when generating training batches, ADMM allows us to tackle the optimization problem holistically. Additionally, it can find a solution for large clinical questionnaire datasets – thousands of participants, tens to hundreds of questions – in about a minute with a laptop CPU, our ballpark for practical usefulness. For smaller-scale problems or validation experiments, we also provide a coordinate descent–based optimization procedure (Appendix B), which relies on inexpensive per-coordinate updates and has a smaller memory footprint, and is primarily intended for modest problem sizes.
@@
$\begin{split}\mathcal{L}_{\rho}(W,Q,Z,\alpha_{Z})=&\;1/2\|\mathcal{M}\odot(M-Z)\|_{F}^{2}+\mathcal{I}_{\mathcal{W}}(W)+\beta\|W\|_{1,1}+\mathcal{I}_{\mathcal{Q}}(Q)+\beta\|Q\|_{1,1}\\ &+\left\langle\alpha_{Z},Z-[W,C]Q^{T}\right\rangle+\rho/2\left\|Z-[W,C]Q^{T}\right\|^{2}_{F}+\mathcal{I}_{\mathcal{Z}}(Z)\end{split}$
@@
Equations 1 and 2 can be further split into row-wise constrained Lasso problems, which can be solved efficiently via an inner-loop ADMM scheme. This is helpful for dealing with large-scale datasets, if memory is limited, and for acceleration if distributed computational resources are available. For small to moderate problem sizes, coordinate descent (CD) is empirically more efficient and simpler to implement. We briefly outline the CD updates for 1; updates for 2 follows analogously. After rescaling and simplifying notation, the subproblem reduces to
$\min_{W\in\mathcal{W}}\;\frac{1}{2}\|V-WH\|_{F}^{2}+\frac{\beta_{W}}{\rho}\|W\|_{1,1},$ (5)
where $V$ and $H$ are fixed. CD updates one entry $W_{ir}$ at a time by solving a one-dimensional quadratic subproblem, yielding the closed-form update
$W_{ir}\leftarrow\Pi_{[0,B_{W}]}\!\left(W_{ir}-\frac{(WHH^{\top}-VH^{\top})_{ir}+\beta_{W}/\rho}{(HH^{\top})_{rr}}\right),$ (6)
where $\Pi_{[0,B_{W}]}(\cdot)$ denotes projection onto the feasible interval. Noted that missing entries in $\mathbf{M}$ do not affect the updates, since the auxiliary variable $Z$ serves as a fully observed surrogate within ADMM. For full derivations and implementation, see Appendix A.
There is a closed form solution for equation 3, since both terms in equation 3 are in Frobenius-norm, so $Z$ can be optimized entry-wise. Specifically, we have the following closed-form solution for $Z^{(i+1)}$:
$Z^{(i+1)}=\underset{[\min(M),\max(M)]}{Proj}\left(\mathcal{M}\odot M+\rho[W^{(i+1)},C](Q^{(i+1)})^{T}-\alpha_{Z}^{(i)}\right)\oslash(\rho\mathbbm{1}+\mathcal{M})$ (7)
[…]
Convergence of the optimization procedure  The convergence hinges on the careful selection of the penalty parameter $\rho$. Informally, imposing the constraint $\rho\geq\sqrt{2}$ on the penalty parameter $\rho$ guarantees monotonicity of the optimization procedure, and that it will converge to a local minimum. Integrating this constraint with the adaptive selection of $\rho$ Xu et al. (2017), we obtain an efficient optimization procedure for ICQF. Formally, this can be stated as the following proposition.
Proposition 3.1 (Non-increasing property).
Assuming $\rho\geq\sqrt{2}$, we have
@@
Proof Sketch: We present a sketch of the proof establishing the monotonicity property of the augmented Lagrangian in Equation 8. Let $\mathbbm{V}^{(i)}=\{W^{(i)},Q^{(i)},Z^{(i)}\}$ and define $R^{(i)}=[W^{(i)},C](Q^{(i)})^{T}$. We examine the difference in the augmented Lagrangian values between two consecutive iterations,
$\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1)},\alpha_{Z}^{(i+1)})-\mathcal{L}_{\rho}(\mathbbm{V}^{(i)},\alpha_{Z}^{(i)})$ $=\underbrace{\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1)},\alpha_{Z}^{(i+1)})-\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1)},\alpha_{Z}^{(i)})}_{(I)}$
$\quad+\underbrace{\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1)},\alpha_{Z}^{(i)})-\mathcal{L}_{\rho}(\mathbbm{V}^{(i)},\alpha_{Z}^{(i)})}_{(II)}.$ (9)
Using the ADMM dual update $\alpha_{Z}^{(i+1)}=\alpha_{Z}^{(i)}+\rho(Z^{(i+1)}-R^{(i+1)})$, term $(I)$ can be written as
$(I)=\langle\alpha_{Z}^{(i+1)}-\alpha_{Z}^{(i)},\,Z^{(i+1)}-R^{(i+1)}\rangle=\frac{1}{\rho}\|\alpha_{Z}^{(i+1)}-\alpha_{Z}^{(i)}\|_{F}^{2}.$ (10)
We further decompose term $(II)$ according to the sequential updates of $Z,Q$, and $W$,
$(II)$ $=\underbrace{\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1)},\alpha_{Z}^{(i)})-\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1,i+1,i)},\alpha_{Z}^{(i)})}_{(\mathcal{A})}+\underbrace{\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1,i+1,i)},\alpha_{Z}^{(i)})-\mathcal{L}_{\rho(\mathbbm{V}^{(i+1,i,i)},\alpha_{Z}^{(i)})}}_{(\mathcal{B})}$
$\quad+\underbrace{\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1,i,i)},\alpha_{Z}^{(i)})-\mathcal{L}_{\rho}(\mathbbm{V}^{(i)},\alpha_{Z}^{(i)})}_{(\mathcal{C})}.$ (11)
The term $(\mathcal{A})$ corresponds to the $Z$-update. By expanding the quadratic terms in the augmented Lagrangian and using the optimality of $Z^{(i+1)}$ for the $Z$-subproblem, which is strongly convex, we obtain
$(\mathcal{A})\leq-\frac{\rho}{2}\|Z^{(i+1)}-Z^{(i)}\|_{F}^{2}.$ (12)
The term $(\mathcal{B})$ captures the effect of the $Q$-update. Each row of $Q^{(i+1)}$ is obtained by solving a constrained Lasso problem. The first-order optimality condition and the convexity of the $\ell_{1}$-norm imply
$(\mathcal{B})\leq-\frac{\rho}{2}\|[W^{(i+1)},C](Q^{(i+1),T}-Q^{(i),T})\|_{F}^{2}.$ (13)
Similarly, the contribution of the $W$-update satisfies
$(\mathcal{C})\leq-\frac{\rho}{2}\|[(W^{(i+1)}-W^{(i)}),C]Q^{(i),T}\|_{F}^{2}.$ (14)
To bound the dual increment, we use the definition of $R^{(i)}$ and obtain
$\|\alpha_{Z}^{(i+1)}-\alpha_{Z}^{(i)}\|_{F}^{2}$ $\leq\|Z^{(i+1)}-Z^{(i)}\|_{F}^{2}+\|[W^{(i+1)},C](Q^{(i+1),T}-Q^{(i),T})\|_{F}^{2}$
$\quad+\|[(W^{(i+1)}-W^{(i)}),C]Q^{(i),T}\|_{F}^{2}.$ (15)
Combining the above bounds yields
@@
$\mathcal{L}_{\rho}(\mathbbm{V}^{(i+1)},\alpha_{Z}^{(i+1)})-\mathcal{L}_{\rho}(\mathbbm{V}^{(i)},\alpha_{Z}^{(i)})\leq\left(\frac{1}{\rho}-\frac{\rho}{2}\right)\cdot(\text{nonnegative terms}),$ (16)
which implies that the augmented Lagrangian is non-increasing whenever $\rho\geq\sqrt{2}$. The full proof of Proposition 8 is given in Appendix C. $\square$
@@
Furthermore, Bjorck et al. (2021) showed that, for non-negative matrix factorizations, if the dimensionality $k$ is the same as that $k^{*}$ of a ground truth solution $(W^{*},Q^{*})$, the error $\|M-WQ^{T}\|_{F}^{2}$ is star-convex towards $(W^{*},Q^{*})$, and the solution is close to a global minimum. However, if $k\neq k^{*}$, the relative error between $W^{*}$ and $W$ increases with $|\sqrt{k/k^{*}}-1|$. Inaccurate estimation of $k^{*}$ thus affects both the interpretability of ($W$, $Q$) and the convergence to global minima. With the bounded constraints imposed on $W$ and $Q$ in ICQF, Popoviciu’s inequality establishes an upper bound for the variances $\sigma_{W}^{2}$ and $\sigma_{Q}^{2}$ of each column in $W$ and $Q$ respectively. To simplify the analysis, we assume equal variances among the columns (generally true). Then we have the following proposition:
Proposition 3.2.
@@
Let $(W^{*},Q^{*})$ be a ground-truth factorization of the given $\mathbf{M}=\mathbf{W}^{*}(\mathbf{Q}^{*})^{T}$, with latent dimension $k^{*}$, where $\mathbf{W}^{*}$ and $\mathbf{Q}^{*}$ are matrix-valued random variables with entries sampled from bounded distributions. Suppose $(\mathbf{W},\mathbf{Q})$ be another factorization with latent dimension $k\neq k^{*}$ that achieves the same expected reconstruction error and the same expected approximation to $\mathbf{M}$, and whose latent coordinates are isotropic in expectation. Then, with high probability,